OpenStudy (anonymous):

help with logs!!! solve for x: 1. 2^2x+2^x-12=0 2. 3^2x+3^x-2=0 3. e^2x-2e^x-3=0

OpenStudy (anonymous):

@5resnz @JCMMusic do you know how to do these?

OpenStudy (anonymous):

No not really sorry.

OpenStudy (anonymous):

is it like this: 1. x = log 3 / log 2 2. x = log(2)/log(3) ≈ 0.6309 3. e^(x) = 3 and e^(x) = -1 x = ln(3) [No x value for e^(x) = -1 since e^(x) is never negative]

OpenStudy (anonymous):

hope it helps but if it doesn't then sry...:)